What is the equation of the line that passes through the points (-1,3), and (0,1)?
Answer by Cromlix(4381)

Hi there,
Gradient = y2 - y1/x2 - x1
Gradient = 1 - 3/0 -(-1)
Gradient = -2/1
Gradient = -2
Using line equation y - b = m(x - a)
m = -2 (a,b) = (-1,3)
y - 3 = -2(x - (-1))
y - 3 = -2(x + 1)
y - 3 = -2x - 2
y = -2x - 2 + 3
y = -2x + 1
Hope this helps :-)